# Ab initio uncertainty quantification in scattering analysis of microscopy

### v2

@@ Abstract
Estimating parameters from data is a fundamental problem in physics, customarily done by minimizing a loss function between a model and observed statistics. In scattering-based analysis, it is common to work in the reciprocal space. Researchers often employ their domain expertise to select a specific range of wavevectors for analysis, a choice that can vary depending on the specific case. We introduce another paradigm that defines a probabilistic generative model from the beginning of data processing and propagates the uncertainty for parameter estimation, termed the ab initio uncertainty quantification (AIUQ). As an illustrative example, we demonstrate this approach with differential dynamic microscopy (DDM) that extracts dynamical information through minimizing a loss function for the squared differences of the Fourier-transformed intensities, at a selected range of wavevectors. We first show that the conventional way of estimation in DDM is equivalent to fitting a temporal variogram in the reciprocal space using a latent factor model as the generative model. Then we derive the maximum marginal likelihood estimator, which optimally weighs the information at all wavevectors, therefore eliminating the need to select the range of wavevectors. Furthermore, we substantially reduce the computational cost of computing the likelihood function without approximation, by utilizing the generalized Schur algorithm for Toeplitz covariances. Simulated studies of a wide range of dynamical systems validate that the AIUQ method improves estimation accuracy and enables model selection with automated analysis. The utility of AIUQ is also demonstrated by three distinct sets of experiments: first in an isotropic Newtonian fluid, pushing limits of optically dense systems compared to multiple particle tracking; next in a system undergoing a sol-gel transition, automating the determination of gelling points and critical exponent; and lastly, in discerning anisotropic diffusive behavior of colloids in a liquid crystal. These studies demonstrate that the new approach does not require manually selecting the wavevector range and enables automated analysis.
@@ I Introduction
In this work, we first introduce a new paradigm that defines a probabilistic generative model from the beginning of data processing. Then we propagate the uncertainty throughout the analysis by integrating out the random quantities, to derive an optimal statistical estimator, such as the maximum marginal likelihood estimator, and subsequently implement it in fast algorithms. Through the scattering analysis, we illustrate that the new estimator automatically weighs information from data at different bases, such as different wavevectors in Fourier-based analysis, thus lifting the barriers for selecting the wavevector range.
@@
We apply our automated and scalable analysis to video microscopy, one of the most ubiquitous tools to access the microscopic realm, including cells, microcapillaries, bacteria, and colloids. Contemporary video microscopy offers not only visual insight but also versatility and power, supporting multiplexing imaging and capturing time sequences of dynamical processes. Approaches to process video microscopy can be broadly defined in two classes: particle tracking and basis decomposition tools. Particle tracking tools start with separating particle intensity profiles from the background for various techniques including fluorescence microscopy [9], total internal reflection fluorescence microscopy [10], and dark field microscopy [11, 12]. Once the particle profiles have been identified in individual frames, the trajectories of particles between consecutive time frames are linked using multiple particle tracking (MPT) algorithms, such as the Crocker-Grier Algorithm [13], also available in MATLAB [14] and Python [15]. ImageJ [16] and Trackmate [17], have been extensively used for analyzing biophysical and cellular processes. Machine learning tools have also been developed for static cellular images, such as the cell profiling tools (e.g. CellProfiler [18]) and segmentation tools (e.g. CellPose [19]). While there are well-established frameworks that explain static and dynamic errors, such as those proposed by [20, 21], MPT algorithms still depend on several user-specified parameters for localization and selection, e.g. the search radius for linking particle trajectories. As a result, particle-based tools can yield unsatisfactory results owing to factors such as irregular shapes, optically dense environments, fluctuations in the size, shape, and fluorescent intensity of moving objects.
@@
On the other hand, basis decomposition tools reconstruct microscopy images through basis functions. One of the representative tools is differential dynamic microscopy, which was introduced in [22, 23]. These seminal works innovatively treat each pixel of the image from a microscopy video as probes in dynamic light scattering [24], as such, they can extract dynamical information of the system by correlating photon counts at distinct time points. In DDM, the squared difference of Fourier-transformed intensity at any two time frames, often referred as the image structure function, is related to the intermediate scattering function [9], which encodes pattern information of dynamical processes. As such, DDM analysis of video microscopy does not require localizing the particles and linking their trajectory. Thus, it is well-suited as a complement to tracking-based tools such as MPT. DDM has been applied to a broad range of soft materials and biological systems, including, for instance, bacteria motility [25, 4], colloidal gels [8], viscoelastic processes [26], active filament dynamics [27], and protein gelation dynamics [28].
@@
However, obstacles remain, preventing DDM from achieving full automation in high-throughput settings. In DDM, one fits the image structure function, which often varies by several orders of magnitude across different wavevectors and furthermore, the data are correlated over different lag time for the same wavevector. As such, this requires specifying the wavevector range or weighing information at wavevectors in a case-by-case manner. The difficulty in selecting wavevector to analyze is common for other scattering approaches, such as dynamic light scattering [24, 29], thus solving it inspires new approaches for a wide range of characterization techniques. Here, we develop a probabilistic latent factor model, which encodes the physics-informed intermediate scattering function as the covariance function of latent factors, and derive the maximum marginal likelihood estimator, therefore removing the need for selecting wavevectors and weighing information in a case-by-case manner. Utilizing fast computational algorithms, such as the generalized Schur algorithm, we also reduce the cost for computing the likelihood function without approximation.
@@
Our contributions are three-fold. First of all, we introduce a new generative model of scattering analysis of microscopy in Sec. III.1. We show that conventional estimation of the parameters in DDM is equivalent to fitting a temporal variogram in the Fourier space using the generative model in Sec. III.2. This connection enhances our understanding of the existing estimator in DDM by integrating intermediate scattering function into the temporal covariance of a latent factor model at each wavevector. We show the conventional practice by fitting the image structure function in DDM is equivalent to minimizing the average of the temporal variogram from our latent factor model. To achieve better efficiency of estimation and remove the need to select wavevectors, we derive the maximum marginal likelihood estimator (MMLE) after integrating out the latent factor processes introduced in Sec. III.3. As information on each wavevector is weighed appropriately by the likelihood, one can utilize all wavevectors instead of tailoring the range for different systems for estimation. Second, directly computing the MMLE is prohibitively slow due to computing inversion and log determinants of a large number of covariance matrices. By evoking the Toeplitz structure of the covariances in the Fourier space, we apply the generalized Schur algorithm [30] to accelerate the computation of the marginal likelihood function without approximation in Sec. IV.1, reducing the computational order from cubic to pseudo-linear (or linear with respective to a log multiplicative constant) scaling to the number of time points, without approximating the likelihood. We show in Sec. IV.2 that computational cost can be further reduced with a principled way of data reduction. Taken together, for a typical microscopy video with $500\times 500$ pixels and $500$ time frames, our approach is more than $10^{5}$ times faster than the direct computation of the likelihood function. Third, the generative model and fast algorithm enable a wide range of applications, such as automated determination of gelation point, discussed in Sec. III.5. Finally, the new probabilistic approach provides uncertainty quantification of the estimation, and the likelihood can be utilized to select physical models by data. We demonstrate the approach by simulation studies and three distinct types of experiments, including optimally dense particles, high-throughput determinant of gelling point, and estimation of anisotropic processes, where the estimation is automated for these applications.
@@ II Background for scattering analysis of microscopy
To start, we consider a system of $M$ particles in a 2D space with $\mathbf{x}_{m}(t)=(x_{m,1}(t),x_{m,2}(t))^{T}$ being the 2D particle location of particle $m$ at time $t$, for $m=1,...,M$. The normalized Fourier-transformed intensity can be written as the sum of the particle positions in the reciprocal space
@@
where $i$ denotes the imaginary unit. Assume the particles do not interact with each other [9]. The intermediate scattering function (ISF), an important function encapsulating the time evolution of particle self-correlation, is characterized by a vector of parameters $\bm{\theta}$, below,
@@
Various processes have a closed-form expression of ISF. A few examples of closed-form ISF are derived in Appendix A and summarized in Table 3. For instance, for Brownian motion (BM) or diffusive processes, the intermediate scattering function is $f_{BM}(\mathbf{q},\Delta t)=\exp(-q^{2}D\Delta t)$, where $q=||\mathbf{q}||$ and here the only parameter $\theta$ of ISF is the diffusion coefficient. By the cumulant theorem [36], the ISF can be approximately characterized by the mean squared displacement (MSD), discussed in Appendix A. Our new approach is generally applicable to all ISFs with a small number of parameters $\bm{\theta}$, and the ISFs are not necessarily approximated by the MSD. We intend to estimate both the parameters $\bm{\theta}$ and system properties such as MSD, which can be related to viscosity, storage and loss modulus through the Generalized Stokes–Einstein Equation (GSER) [37].
@@
$D(\mathbf{q},\Delta t)=A(\mathbf{q})(1-f_{\bm{\theta}}(\mathbf{q},\Delta t))+B,$ (3)
where $A(\mathbf{q})$ is the real-valued scalar of amplitude parameter for wavevector $\mathbf{q}$, $f_{\bm{\theta}}$ is the ISF defined in Eq. (2), $B$ denotes a random noise with a mean value $\bar{B}$. For an isotropic process in a square field of view with $N$ pixels, we denote an index set $\mathcal{S}_{j}=\{(j^{\prime}_{1},j^{\prime}_{2}):q^{2}_{j^{\prime}_{1},1}+q^{2}_{j^{\prime}_{2},2}=q^{2}_{j}\}$ for $j=1,...,J$, which contains the indices of the $j$th ‘ring’ of the Fourier-transformed quantity with amplitude $q_{j}=\frac{2\pi j}{\Delta x_{min}\sqrt{N}}$, with $\Delta x_{min}$ being the pixel size, i.e. the length of a pixel in one coordinate, and $N$ is the number of pixels in one frame. In total, there are J rings of Fourier-transformed intensities, leading to $J$ distinct ISFs for isotropic processes. Furthermore, assuming we have n time frames, denote $\mathbf{D}$ as the $J\times(n-1)$ matrix with $(j,k)$th term being the observed image structure function $D(q_{j},\Delta t_{k})=\langle\Delta{\hat{y}}(\mathbf{q}_{\mathbf{j}^{\prime}},t,\Delta t_{k})^{2}\rangle$, where now the ensemble is over both time $t$ and indices within each ring $\mathbf{j}^{\prime}\in\mathcal{S}_{j}$. Let $\mathbf{D}_{m}$ denote the model output of $\mathbf{D}$, where the $(j,k)$th entry of $\mathbf{D}_{m}$ is $D_{m}(q_{j},\Delta t_{k})=A_{j}(1-f_{\bm{\theta}}(q_{j},\Delta t_{k}))+\bar{B}$ with $\bar{B}$ denoting the mean for the random quantity $B$. In DDM, the parameters are often estimated by minimizing a loss function between the observed and modeled image structure functions
@@
with $\mathbf{A}_{est}=[A_{est,1},...,A_{est,J}]^{T}$ being a J-vector of amplitudes. A typical choice of the loss function is either the $L_{2}$ or $L_{1}$ loss. As the parameter space has a high dimension, the ISF is often fit separately for each wavevector from a selected range, distinct in each application [22, 4, 39, 40], and then estimators at different wavevectors are averaged to obtain $\bm{\theta}$. Some variants of DDM [41, 42, 43] use different pre-specified estimators for $\bar{B}_{est}$ and an unbiased estimator of $A_{est,j}=2\langle\hat{y}(\mathbf{q}_{\mathbf{j}^{\prime}},t)^{2}\rangle_{\mathbf{j}^{\prime}\in\mathcal{S}_{j},t}-\bar{B}_{est}$ to estimate $A_{j}$, leaving $\bm{\theta}$ in the ISF the only parameters to be numerically optimized. A summary of the estimators $\bar{B}_{est}$ is introduced in [44]. In [26, 44], the ISF is approximated by MSD at wavevector, i.e. $f_{\bm{\theta}}(q,\Delta t)\approx\exp\left(-q^{2}\langle\Delta x^{2}(\Delta t)\rangle/4\right)$, and Eq. (3) is directly inverted at each $\Delta t$ to obtain the estimator of MSD $\langle x^{2}(\Delta t)\rangle$. In [45], the MSD is obtained through iterative optimization to reduce numerical instability.
However, estimation by directly fitting or inverting the image structure function can depend on the range of wavevector selected. This is illustrated by simulating a simple diffusion process, shown in Figure 9 in Appendix B. Almost all existing approaches in DDM fit the image structure function (Eq. (3)), relying on selecting a subset of wavevectors to analyze, whereas selecting and reweighing the information at different wavevectors can be hard for a new system. This difficulty arises from the substantial variations in amplitudes $A_{j}$ and the correlation of the image structure function at different lag times. A principled way to properly aggregate information at different wavevectors can unlock the tremendous potential for DDM to attain complete automation in this process, yet this has not been fully realized so far.

### v4

@@ Abstract
Estimating parameters from data is a fundamental problem in physics, customarily done by minimizing a loss function between a model and observed statistics. In scattering-based analysis, it is common to work in the reciprocal space. Researchers often employ their domain expertise to select a specific range of wave vectors for analysis, a choice that can vary depending on the specific case. We introduce another paradigm that defines a probabilistic generative model from the beginning of data processing and propagates the uncertainty for parameter estimation, termed the ab initio uncertainty quantification (AIUQ). As an illustrative example, we demonstrate this approach with differential dynamic microscopy (DDM) that extracts dynamical information through minimizing a loss function for the squared differences of the Fourier-transformed intensities, at a selected range of wave vectors. We first show that the conventional way of estimation in DDM is equivalent to fitting a temporal variogram in the reciprocal space using a latent factor model as the generative model. Then we derive the maximum marginal likelihood estimator, which optimally weighs the information at all wave vectors, therefore eliminating the need to select the range of wave vectors. Furthermore, we substantially reduce the computational cost of computing the likelihood function without approximation, by utilizing the generalized Schur algorithm for Toeplitz covariances. Simulated studies of a wide range of dynamical systems validate that the AIUQ method improves estimation accuracy and enables model selection with automated analysis. The utility of AIUQ is also demonstrated by three distinct sets of experiments: first in an isotropic Newtonian fluid, pushing limits of optically dense systems compared to multiple particle tracking; next in a system undergoing a sol-gel transition, automating the determination of gelling points and critical exponent; and lastly, in discerning anisotropic diffusive behavior of colloids in a liquid crystal. These studies demonstrate that the new approach does not require manually selecting the wave-vector range and enables automated analysis.
@@ I Introduction
In this work, we first introduce a new paradigm that defines a probabilistic generative model from the beginning of data processing. Then we propagate the uncertainty throughout the analysis by integrating out the random quantities, to derive an optimal statistical estimator, such as the maximum marginal likelihood estimator, and subsequently implement it in fast algorithms. Through the scattering analysis, we illustrate that the new estimator automatically weighs information from data at different bases, such as different wave vectors in Fourier-based analysis, thus lifting the barriers for selecting the wave-vector range.
@@
We apply our automated and scalable analysis to video microscopy, one of the most ubiquitous tools to access the microscopic realm, including cells, bacteria, and colloids. Contemporary video microscopy offers not only visual insight but also versatility and power, supporting multiplexing imaging and capturing time sequences of dynamical processes. Approaches to process video microscopy can be broadly defined in two classes: particle tracking and basis decomposition tools. Particle tracking tools start with separating particle intensity profiles from the background for various techniques including fluorescence microscopy [9], total internal reflection fluorescence microscopy [10], and dark field microscopy [11, 12]. Once the particle profiles have been identified in individual frames, the trajectories of particles between consecutive time frames are linked using multiple particle tracking (MPT) algorithms, such as the Crocker-Grier Algorithm [13], also available in MATLAB [14] and Python [15]. ImageJ [16] and Trackmate [17], have been extensively used for analyzing biophysical and cellular processes. Machine learning tools have also been developed for static cellular images, such as the cell profiling tools (e.g. CellProfiler [18]) and segmentation tools (e.g. CellPose [19]). While there are well-established frameworks that explain static and dynamic errors, such as those proposed by [20, 21], MPT algorithms still depend on several user-specified parameters for localization and selection, e.g. the search radius for linking particle trajectories. As a result, particle-based tools can yield unsatisfactory results owing to factors such as irregular shapes, optically dense environments, fluctuations in the size, shape, and fluorescent intensity of moving objects.
@@
On the other hand, basis decomposition tools reconstruct microscopy images through basis functions. A class of basis decomposition methods termed Digital Fourier Microscopy (DFM) performs temporal analysis of the Fourier-transformed microscopy image sequences to extract process dynamics [22]. This approach originates from scattering, which collects signals in the far field and computes the ensemble averages over many scatterers. It has been shown that near-field microscopy images encode the information on the correlation properties of dynamical processes analogous to scattering [23, 24]. Among them, a representative method is differential dynamic microscopy, which was introduced in [23, 24]. These seminal works innovatively treat each pixel in the Fourier space transformed from a microscopy video as probes in dynamic light scattering [25], as such, they can extract dynamical information of the system by correlating photon counts at distinct time points. In DDM, the squared difference of Fourier-transformed intensity at any pairs of frames, often referred to as the image structure function, is related to the intermediate scattering function [9]. DDM analysis of video microscopy does not require localizing the particles and linking their trajectory, provides access to high-quality analysis of the dynamics and requires no specialized imaging source or setup. Thus, it serves as a complement to tracking-based tools such as MPT. DDM has been applied to a broad range of soft materials and biological systems, including, for instance, bacteria motility [26, 4], colloidal gels [8], viscoelastic processes [27], active filament dynamics [28], and protein gelation dynamics [29].
@@
However, obstacles remain, preventing DDM from achieving full automation in high-throughput settings. In DDM, one fits the image structure function, which often varies by several orders of magnitude across different wave vectors and furthermore, the data are correlated over different lag time for the same wave vector. As such, this requires specifying the wave-vector range or weighing information at wave vectors in a case-by-case manner. The difficulty in selecting wave vector to analyze is common for other scattering approaches, such as dynamic light scattering [25, 30]. Thus solving it inspires new approaches for a wide range of characterization techniques. Here, we develop a probabilistic latent factor model, which encodes the physics-informed intermediate scattering function as the covariance function of latent factors, and derive the maximum marginal likelihood estimator, therefore removing the need for selecting wave vectors and weighing information in fitting the image structure function. Utilizing fast computational algorithms, such as the generalized Schur algorithm, we also reduce the cost for computing the likelihood function without approximation.
@@
Our contributions are threefold. First of all, we introduce a new probabilistic generative model of scattering analysis of microscopy in Sec. III.1. We show that conventional estimation of the parameters in DDM is equivalent to fitting a temporal variogram in the Fourier space using the generative model in Sec. III.2. This connection enhances our understanding of the existing estimator in DDM by integrating intermediate scattering function into the temporal covariance of a latent factor model at each wave vector. We show the conventional practice by fitting the image structure function in DDM is equivalent to minimizing the average of the temporal variogram from our latent factor model. To achieve better efficiency of estimation and remove the need to select wave vectors, we derive the maximum marginal likelihood estimator (MMLE) after integrating out the latent factor processes introduced in Sec. III.3. As information on each wave vector is weighed appropriately by the likelihood, one can utilize all wave vectors instead of tailoring the range for different systems for estimation. Second, directly computing the MMLE is prohibitively slow due to computing inversion and log determinants of a large number of covariance matrices. By evoking the Toeplitz structure of the covariances in the Fourier space, we apply the generalized Schur algorithm [31] to accelerate the computation of the marginal likelihood function without approximation in Sec. IV.1, reducing the computational order from cubic to pseudo-linear (or linear with respective to a log multiplicative constant) scaling to the number of time points, without approximating the likelihood. We show in Sec. IV.2 that computational cost can be further reduced with a principled way of data reduction. Taken together, for a typical microscopy video with $500\times 500$ pixels and $500$ time frames, our approach is more than $10^{5}$ times faster than the direct computation of the likelihood function. Third, the generative model and fast algorithm enable a wide range of applications, such as automated determination of gelation point, discussed in Sec. III.5. Finally, the new probabilistic approach provides uncertainty quantification of the estimation, and the likelihood can be utilized to select physical models by data. We demonstrate the approach by simulation studies and three distinct types of experiments, including optically dense particles, high-throughput determinant of gelling point, and estimation of anisotropic processes, where the estimation is automated for these applications.
@@ II Background for scattering analysis of microscopy
To start, we consider a system of $M$ particles in a two-dimensional (2D) space with $\mathbf{x}_{m}(t)=(x_{m,1}(t),x_{m,2}(t))^{T}$ being the 2D particle location of particle $m$ at time $t$, for $m=1,...,M$. The normalized Fourier-transformed intensity can be written as the sum of the particle positions in the reciprocal space
@@
where $i$ denotes the imaginary unit and $\mathbf{q}$ is a 2D Fourier basis set or a wave vector. Assume the particles do not interact with each other [9]. The intermediate scattering function (ISF), an important function encapsulating the time evolution of particle self-correlation, is characterized by a vector of parameters $\bm{\theta}$, below,
@@
Various processes have a closed-form expression of ISF. A few examples of closed-form ISF are derived in Appendix A and summarized in Table 3. For instance, for Brownian motion (BM) or diffusive processes, the intermediate scattering function is $f_{BM}(\mathbf{q},\Delta t)=\exp(-q^{2}\theta\Delta t)$, where $q=||\mathbf{q}||$ and here the only parameter $\theta$ of ISF is the diffusion coefficient. By the cumulant theorem [37], the ISF can be approximately characterized by the mean squared displacement (MSD), discussed in Appendix A. Our new approach is generally applicable to all ISFs with a vector of parameters $\bm{\theta}$, and the ISFs are not necessarily approximated by the MSD. We intend to estimate both the parameters $\bm{\theta}$ and system properties such as MSD, which can be related to viscosity, storage and loss modulus through the Generalized Stokes–Einstein Equation (GSER) [38].
@@
$D(\mathbf{q},\Delta t)=A(\mathbf{q})(1-f_{\bm{\theta}}(\mathbf{q},\Delta t))+\bar{B},$ (3)
where $A(\mathbf{q})$ is the real-valued scalar of amplitude parameter for wave vector $\mathbf{q}$, $f_{\bm{\theta}}$ is the ISF defined in Eq. (2), $\bar{B}$ denotes with mean value of the noise term. For an isotropic process in a square field of view with $N$ pixels, we denote an index set $\mathcal{S}_{j}=\{(j^{\prime}_{1},j^{\prime}_{2}):q^{2}_{j^{\prime}_{1},1}+q^{2}_{j^{\prime}_{2},2}=q^{2}_{j}\}$ for $j=1,...,J$, which contains the indices of the $j$th ‘ring’ of the Fourier-transformed quantity with amplitude $q_{j}=\frac{2\pi j}{\Delta x_{min}\sqrt{N}}$, with $\Delta x_{min}$ being the pixel size, i.e. the length of a pixel in one coordinate, and $N$ is the number of pixels in one frame. In total, there are $J$ rings of Fourier-transformed intensities, leading to $J$ distinct ISFs for isotropic processes. Furthermore, assuming we have n time frames, denote $\mathbf{D}$ as the $J\times(n-1)$ matrix with $(j,k)$th term being the observed image structure function $D(q_{j},\Delta t_{k})=\langle\Delta{\hat{y}}(\mathbf{q}_{\mathbf{j}^{\prime}},t,\Delta t_{k})^{2}\rangle$, where now the ensemble is over both time $t$ and indices within each ring $\mathbf{j}^{\prime}=(j^{\prime}_{1},j^{\prime}_{2})\in\mathcal{S}_{j}$. Let $\mathbf{D}_{m}$ denote the model output of $\mathbf{D}$, where the $(j,k)$th entry of $\mathbf{D}_{m}$ is $D_{m}(q_{j},\Delta t_{k})=A_{j}(1-f_{\bm{\theta}}(q_{j},\Delta t_{k}))+\bar{B}$ with $\bar{B}$ denoting the mean for the random quantity $B$. In DDM, the parameters are often estimated by minimizing a loss function between the observed and modeled image structure functions
@@
with $\mathbf{A}_{est}=[A_{est,1},...,A_{est,J}]^{T}$ being a J-vector of amplitudes. A typical choice of the loss function is either the $L_{2}$ or $L_{1}$ loss. As the parameter space has a high dimension, one often fits the model separately for each wave vector from a selected range, distinct in each application [23, 4, 40, 41], and then estimators at different wave vectors are averaged to obtain $\bm{\theta}$. Some variants of DDM [42, 43, 44] use different pre-specified estimators for $\bar{B}_{est}$ and an unbiased estimator of $A_{est,j}=2\langle\hat{y}(\mathbf{q}_{\mathbf{j}^{\prime}},t)^{2}\rangle_{\mathbf{j}^{\prime}\in\mathcal{S}_{j},t}-\bar{B}_{est}$ to estimate $A_{j}$, leaving $\bm{\theta}$ in the ISF the only parameters to be numerically optimized. A summary of the estimators $\bar{B}_{est}$ is introduced in [45]. In [27, 45], the ISF is approximated by MSD, i.e. $f_{\bm{\theta}}(q,\Delta t)\approx\exp\left(-q^{2}\langle\Delta x^{2}(\Delta t)\rangle/4\right)$, and Eq. (3) is directly inverted at each $\Delta t$ to obtain the estimator of MSD $\langle x^{2}(\Delta t)\rangle$. In [46], the MSD is obtained through iterative optimization to reduce numerical instability.
However, estimation by directly fitting or inverting the image structure function can depend on the range of wave vector selected. This is illustrated by simulating a simple diffusion process, shown in Figure 9 in Appendix B. Almost all existing approaches in DDM fit the image structure function in Eq. (3), relying on selecting a subset of wave vectors to analyze, whereas selecting and reweighing the information at different wave vectors can be hard for a new system. This difficulty arises from the substantial variations in amplitudes $A_{j}$ and the correlation of the image structure function at different lag times. A principled way to properly aggregate information at different wave vectors can unlock the tremendous potential for DDM to attain complete automation in this process, yet this has not been fully realized so far.